OpenStudy (anonymous):

How do I find the delta of lim x^3=8 x->2 for an epsilon=0.001

OpenStudy (unklerhaukus):

what does epsilon represent

OpenStudy (unklerhaukus):

what does; epsilon, delta represent

OpenStudy (unklerhaukus):

\[\lim\limits_{x\rightarrow 2}x^3=8\]\[\epsilon =0.001\]\[\delta=?\]

OpenStudy (anonymous):

I know that I have to do \[\left| x-8 \right|>\epsilon\] if \[0>\left| x-2 \right|>\delta \]

OpenStudy (unklerhaukus):

how can the absolute value of something be less than zero?

OpenStudy (unklerhaukus):

\[0<|x−2|<δ \]

OpenStudy (anonymous):

u have \[|x^3-8|<\epsilon=0.001\]and u want to find delta with \[|x-2|<\delta\]which implies that delta

OpenStudy (anonymous):

sorry *implies that epsilon

OpenStudy (anonymous):

hint \[-0.001+8<x^3<0.001+8\]apply \(\sqrt[3]{}\)